# arXiv:2603.00321v1[hep-th] 27 Feb 2026

### v1

@@ V Conclusion
In summary, we have generalized the universal relationship between the CFT coefficient $C_{T}$ and the central charge $c$ to arbitrary even dimensions. This highlights the nontrivial connection between conformal field theories in flat space and the trace anomaly arising from gravitational coupling. Notably, several other contexts have established a similar connection to $C_{T}$. These include
@@
A direct mapping between the above quantities and the Weyl anomaly coefficients appears to be a compelling possibility. Our results may find applications in several contexts, such as $O(N)$ vector models and supersymmetric extensions. In the interest of universality, it would also be instructive to consider the energy flux parameters $t_{2}$ and $t_{4}$. These parameters govern the three-point function of the energy-momentum tensor and are expected to relate to Weyl-cubed terms in the trace anomaly in six or more dimensions.
@@ VI Acknowlegments
We are grateful to G.Anastasiou, J.S.Case, F.Diaz, M.J.Duff, A.A.Tseytlin and O.Zanusso for useful discussions, clarifications, and for pointing out relevant references. DED would also like to thank the faculty and staff of the Science Department at Valencia College for their warm welcome and support. This work was partially funded through FONDECYT-Chile 1220335. RA wishes to thank professors J. Gomis (U. Barcelona), E. Bergshoeff (U. Groningen), M. Romo and Hamed Adami (SIMIS, Shanghai), and E. Joung (U. Kyung Hee, Seoul) for their kind hospitality during the completion of this work.

### v2

@@ V Conclusion
In summary, we have generalized the universal relationship between the CFT coefficient $C_{T}$ and the central charge $c$ to arbitrary even dimensions. This highlights the nontrivial connection between conformal field theories in flat space and the trace anomaly arising from gravitational coupling. We have presented three different derivations, a holographic one and two genuine CFT ones. Perhaps the most succinct (and surprising) way to state the universal relation is in terms of the coefficient $C_{TT}$ of the TT-correlator in momentum space (as flat as it gets) and the critical Q-curvature $Q_{d}$ (as curved as it is)
$\boxed{\langle T\rangle\,=\,2\,C_{TT}\cdot Q_{d}+lot+ttd.}$ (64)
This effectively extends to higher-even dimensions previously known results [1, 31, 27, 37] and, at the same time, contains the Weyl-completion of the term quadratic in the Weyl tensor (cf. [55], eqn.3.17 and footnote 16).
Notably, several other contexts have established a similar connection to $C_{T}$. These include
@@
A direct mapping between the above quantities and the Weyl anomaly coefficients appears compelling. In the interest of universality, it would also be instructive to consider the energy flux parameters $t_{2}$ and $t_{4}$. These parameters govern the three-point function of the energy-momentum tensor and are expected to relate to Weyl-cubed terms in the trace anomaly in six or more dimensions.
@@ VI Acknowlegments
We are grateful to G.Anastasiou, J.S.Case, F.Diaz, M.J.Duff, P.McFadden, A.A.Tseytlin and O.Zanusso for useful discussions, clarifications, and for pointing out relevant references. DED would also like to thank the faculty and staff of the Science Department at Valencia College for their warm welcome and support. This work was partially funded through FONDECYT-Chile 1220335. RA wishes to thank professors J. Gomis (U. Barcelona), E. Bergshoeff (U. Groningen), M. Romo and Hamed Adami (SIMIS, Shanghai), and E. Joung (U. Kyung Hee, Seoul) for their kind hospitality during the completion of this work.
Note added: After completion of the first version of this work, we were informed by P. McFadden that in [55], the authors carried out the equivalent to our standard CFT derivation in momentum space and were able to relate their $C_{TT}$ to the coefficient of the $W(\nabla^{2})^{d/2-2}W$ term in the trace anomaly. Their result turns out to be correct in $4D$, but the appropriate extension to higher even dimensions requires an additional dimensional factor (d2)/2(d3)(d-2)/2(d-3) that escaped notice. However, we find their relation very instructive and, consequently, we have casted our own result in terms of $C_{TT}$ and absorbed the dimensional factor into the critical Q-curvature $Q_{d}$, giving a new turn to the universal relation.